A dealer mixes a lbs. of nuts worth b cents per pound with c lbs. of nuts worth d cents per pound. At what price should he sell a pound of the MIXTURE IF HE WISHES TO MAKE A PROFIT OF 10 CENTS PER POUND?
Answer by mananth(16946)

let x be the cost of the mixture
(x+10) will be the price
a*b + c*d = x*(a+c) + 10(a+c)
(ab+cd)= (a+c)(x+10)
(ab+cd)/(a+c) = x+10
[(ab+cd)/(a+c)] = x+10